OpenStudy (anonymous):

Find the vertices and foci of the hyperbola with equation: (x+1)^2/16 - (y+5)^2/9 = 1

OpenStudy (amistre64):

is there a solution when x part = 0? is there a solution when y part = 0?

OpenStudy (amistre64):

might be better stated as: (x+1) = 0 (y+5) = 0

OpenStudy (anonymous):

Okay I see... It just really wants me to write Vertices: ( x, y) and (x, y) then foci: (x,y) and (x, y) @amistre64

OpenStudy (amistre64):

yes, but they have gone and moved this off the origin .... assume the x parts are 0\[-\frac{[~y~]^2}{9}=1\] has no "real" solution so the vertex will have to be when the y parts are zero; or when y = -5

OpenStudy (amistre64):

by zeriong out the y parts we are left with\[\frac{(x+1)^2}{16}=1\] what solutions do we have for this?

OpenStudy (amistre64):

the foci will be relatively simple to determine

OpenStudy (anonymous):

I feel like I'm kind of wasting your time, Amistre... Idk how to find the solutions for that there :(

OpenStudy (amistre64):

at the moment we have Vertices: ( x,-5) and (x, -5) foci: (x,-5) and (x, -5)

OpenStudy (amistre64):

would you agree that anything divided by itself is equal to 1? except for 0/0 of course

OpenStudy (anonymous):

yes. and anything divided by the negative version or negative by positive version is -1

OpenStudy (amistre64):

\[\frac{(x+1)^2}{16}=1\] \[\frac{(x+1)^2}{16}=\frac{16}{16}\] \[(x+1)^2=16\]

OpenStudy (amistre64):

what squared, equals 16?

OpenStudy (anonymous):

4

OpenStudy (amistre64):

4 is one solution, but isnt (-4)^2 = 16 true as well?

OpenStudy (anonymous):

yes so +- 4

OpenStudy (amistre64):

good, then we know that x+1 = 4 or x+1 = -4 what solutions do we have for x now?

OpenStudy (anonymous):

x = 3 or -5

OpenStudy (amistre64):

good, then we know the vertexes verts: (-5,-5) and (3,-5) just have to determine the foci

OpenStudy (anonymous):

okay lets do this

OpenStudy (amistre64):

add the under parts together .... and sqrt the results, we will need this for a measurement

OpenStudy (amistre64):

the general hyperB eq is \[\frac{x^2}{a^2}-\frac{y^2}{b^2}=1\] the foci have a measurement from the center that is equal to \(\sqrt{a^2+b^2}\)

OpenStudy (anonymous):

okay but we never wrote x^2 and y^2 over anything

OpenStudy (amistre64):